### v3

# MERaLiON: Enhancing Cross-Lingual Understanding of Chinese, Indonesian, Malay and Singlish in Language Models

Abstract
Multilingual large language models (MLLMs) have shown impressive capabilities across a variety of languages. However, efficacy can differ greatly between different language families, especially for those with limited linguistic resources. This report presents MERaLiON-TextLLM, a series of open-source language models specifically tailored to improve understanding and generation in Chinese, Indonesian, Malay, and Singlish. The initial released model is built on Llama-3-8B-Base and refined through a meticulously crafted process of continued pre-training and weight merging. Our approach achieves performance improvements across benchmarks in these languages, exceeding the capabilities of the official Llama-3 models. We provide the model checkpoints as a resource to support further research and development in cross-lingual language understanding.111 Please cite this report as authored by MERaLiON team.
Corresponding: {huangx2, vangani_tarun_kumar}@i2r.a-star.edu.sg
Available at: https://huggingface.co/MERaLiON/LLaMA-3-MERaLiON-8B-Instruct
1 Introduction
Our first released model is named LLaMA-3-MERaLiON-8B-Instruct. LLaMA-3-MERaLiON-8B-Instruct has been extensively pre-trained on English, Chinese, and Indonesian, building upon LLaMA-3.1-8B-Base Dubey et al. (2024), with a primary emphasis on enhancing its understanding and generation capabilities in Chinese and Indonesian. Leveraging innovative corpus mixing strategies tailored to multilingual regional datasets, we carefully diversified the training materials using domain classification, hyperparameter optimization, and strategic replay techniques. These methods are specifically designed to prevent catastrophic forgetting, enabling the model to retain previously acquired knowledge while significantly improving its ability to produce high-quality, contextually accurate responses in Southeast Asian languages.
2 Pre-Training
LLaMA-3-MERaLiON-8B-Instruct training was conducted using the MaxText AI-Hypercomputer (2024) platform, utilizing both NVIDIA H100 GPUs and TPU v4-128 chips. Specifically, we utilized 64 H100 GPUs, achieving approximately 400 TFLOPS per GPU, and TPU v4-128 configurations, attaining around 168 TFLOPS per TPU chip.
These performance metrics were achieved through optimized sharding, checkpoint strategies, and the selection of optimal batch sizes, to ensure efficient and effective model training. We conducted multiple runs to achieve optimal performance and sampling strategies.
The pre-training data was allocated across three primary languages: English, Indonesian, and Chinese. Specifically, 38 billion tokens were allocated to English, 45 billion to Indonesian, and 42 billion to Chinese. This allocation is designed to ensure robust performance across all three languages, catering to a variety of linguistic and cultural contexts.
Figure 1 illustrates the token distribution in a pie chart format for clear visualization. English accounted for 30.4% of the total tokens, Indonesian for 36.0%, and Chinese for 33.6%. The allocation highlights the model’s multilingual focus, ensuring no language dominates the training set, thereby fostering balanced performance.
[…]
Figure 1: Distribution of Tokens
3 Instruction Tuning and Model Merging
Table 1: Performance comparison of the merged model (MERaLiON-LLaMA-3.1-8B-Instruct) against baseline and instruction-tuned models on Cross-MMLU and Cross-LogiQA benchmarks.
Model Cross-MMLU Cross-LogiQA English Chinese Indonesian English Chinese Indonesian LLaMA-3-MERaLiON-8B-Instruct 0.85 0.69 0.71 0.591 0.528 0.494 Meta-Llama-3.1-8B-Instruct 0.82 0.63 0.66 0.585 0.585 0.455 Instruction-Tuned Llama-3.1-8B 0.73 0.60 0.58 0.563 0.523 0.466
We present updated experimental results for our latest variant of the Southeast Asian language model, focusing on three representative languages: English, Chinese, and Indonesian. Our work began with the construction of a multilingual instruction-response corpus containing approximately 3M pairs. Using this dataset, we performed instruction tuning on the Llama-3.1-8B base model to evaluate potential performance gains. However, our initial experiments revealed that this approach alone did not surpass the robust baseline performance of the Llama-3.1-8B-Instruct model. Specifically, the instruction-tuned model underperformed the original Llama-3.1-8B-Instruct across the three target languages in the Cross-MMLU and Cross-LogiQA datasets Wang et al. (2024).
To address these limitations, we explored model merging techniques aimed at enhancing instruction-following capabilities without requiring extensive additional tuning. By carefully merging the weight sets of Llama-3.1-8B-Base and Llama-3.1-8B-Instruct, we leveraged the strong instruction-following performance of the instructed variant while integrating domain-specific knowledge from our custom dataset. As illustrated in Table 1, the resulting "MERaLiON-Llama-3.1-8B-Instruct" model consistently outperformed both the baseline Llama-3.1-8B-Instruct and our standalone instruction-tuned variant across English, Chinese, and Indonesian test sets.
These results underscore the effectiveness of merging pretrained instructed weights with domain-specific models to achieve enhanced instruction-following capabilities. The MERaLiON-Llama-3.1-8B-Instruct model demonstrates improved performance in all evaluated languages, confirming the viability of this approach. This hybrid methodology offers a promising avenue for the development of resource-efficient, multilingual, and domain-adaptive instruction-following models without necessitating extensive iterative tuning on large-scale, specialized corpora.
4 Benchmarks and Evaluation
Table 2: Cross-MMLU
Model Series Model English Chinese Indonesian Malay Avg LLaMA3 LLaMA-3-MERaLiON-8B-Instruct 0.847 0.693 0.713 0.613 0.717 Meta-Llama-3.1-8B-Instruct 0.82 0.633 0.66 0.647 0.690 Llama3-8B-SEA-LION-v2.1-Instruct Singapore (2024) 0.753 0.667 0.693 0.64 0.688 Meta-Llama-3-8B-Instruct 0.767 0.653 0.573 0.573 0.642 Others GPT4o-0513 0.927 0.887 0.88 0.907 0.900 Gemma-2-9B-IT Team et al. (2024) 0.84 0.793 0.78 0.747 0.790 Gemma2-9B-SEA-LION-v3-Instruct Singapore (2024) 0.847 0.787 0.793 0.733 0.790 Qwen2.5-7B-Instruct Yang et al. (2024) 0.847 0.84 0.753 0.713 0.788 SeaLLMs-v3-7B-Chat Zhang et al. (2024) 0.833 0.727 0.74 0.687 0.747
Table 3: Cross-LogiQA
Model Series Model English Chinese Indonesian Malay Avg LLaMA3 Meta-Llama-3.1-8B-Instruct 0.585 0.585 0.455 0.523 0.537 LLaMA-3-MERaLiON-8B-Instruct 0.591 0.528 0.494 0.489 0.526 Meta-Llama-3-8B-Instruct 0.602 0.523 0.438 0.483 0.512 Llama3-8B-SEA-LION-v2.1-Instruct 0.528 0.517 0.403 0.443 0.473 Others Qwen2.5-7B-Instruct 0.693 0.71 0.631 0.534 0.642 Gemma-2-9B-IT 0.659 0.636 0.585 0.602 0.621 Gemma2-9B-SEA-LION-v3-Instruct 0.636 0.642 0.557 0.551 0.597 SeaLLMs-v3-7B-Chat 0.568 0.585 0.494 0.517 0.541
We conducted comprehensive evaluations using several benchmarks to assess the multilingual and instruction-following performance of MERaLiON. Key benchmarks include:
  • Cross-MMLU Wang et al. (2024) : A subset of the MMLU dataset translated into multiple languages, including English, Chinese, Indonesian, and Malay. It aims to measure the model’s ability to handle general knowledge queries across these diverse linguistic contexts.
  • Cross-LogiQA Wang et al. (2024) : Building on the original LogiQA dataset, Cross-LogiQA introduces multilingual versions of logical reasoning tasks. It provides parallel question sets in English, Chinese, and Indonesian that are designed to maintain logical equivalence across these languages.
  • IndoMMLU Koto et al. (2023) : A benchmark designed to assess general knowledge and language understanding of large language models for Indonesian, particularly for domains such as law, medicine, and social science Koto et al. (2023).
  • CN-Eval: A selected subset of C-Eval Huang et al. (2023) and CMMLU Li et al. (2024) curated to specifically assess a model’s knowledge about the Chinese language, culture and socio-political context. This subset provides a focused metric for assessing LLM knowledge of China.
4.1 Results
The results of our evaluation highlight several strengths of MERaLiON:
Table 4: IndoMMLU
Model Series Model Accuracy LLaMA Series LLaMA-3-MERaLiON-8B-Instruct 0.576 Llama3-8B-SEA-LION-v2.1-Instruct 0.560 Meta-Llama-3.1-8B-Instruct 0.548 Meta-Llama-3-8B-Instruct 0.521 Others GPT4o-0513 0.760 Gemma2-9B-SEA-LION-v3-Instruct 0.626 Gemma-2-9B-IT 0.621 Qwen2.5-7B-Instruct 0.582 SeaLLMs-v3-7B-Chat 0.541
Table 5: CN-Eval
Model Series Model Accuracy LLaMA Series LLaMA-3-MERaLiON-8B-Instruct 0.514 Llama3-8B-SEA-LION-v2.1-Instruct 0.505 Llama3-8B-SEA-LION-v2-Instruct 0.495 Meta-Llama-3-8B-Instruct 0.467 Meta-Llama-3.1-8B-Instruct 0.457 Others Qwen2-7B-Instruct 0.829 GPT4o-0513 0.810 Qwen2.5-7B-Instruct 0.800 Gemma2-9B-SEA-LION-v3-Instruct 0.590 Gemma-2-9B-IT 0.581
  • On Cross-MMLU and Cross-LogiQA, as shown in Table 2 and Table 3, MERaLiON outperforms the baseline Meta-Llama-3.1-8B-Instruct model on Cross-MMLU, demonstrating superior general knowledge coverage in Chinese and Indonesian, while maintaining a strong, balanced performance in English. Similarly, MERaLiON improves logical reasoning in Indonesian, indicating that its continued pre-training and model merging strategies effectively enhance both factual knowledge and reasoning capabilities.
  • On IndoMMLU, as reported in Table 4, MERaLiON significantly outperforms the baseline Llama-3.1-8B-Instruct model, achieving 0.576 accuracy versus 0.548, highlighting its improved understanding of Indonesian language and domain-specific nuances.
  • For CN-Eval, MERaLiON achieves 0.514 accuracy compared to 0.457 for Llama-3.1-8B-Instruct, as shown in Table 5. This result demonstrates MERaLiON’s efficacy in retaining and enhancing knowledge related to China.
These benchmarks demonstrate the superior capability of MERaLiON-TextLLM to handle multilingual tasks and deliver improved cross-lingual understanding. Our training approach ensures the model is well-suited for diverse language tasks.
5 Conclusion and Future Work
The LLaMA-3-MERaLiON-8B-Instruct model improves multilingual NLP for Indonesian and Chinese, by addressing cross-lingual challenges with effective pretraining and model merging.
Evaluation results demonstrate MERaLiON-TextLLM’s strengths, such as better reasoning and question-answering on Cross-MMLU and Cross-LogiQA through balanced multilingual training, strong accuracy on IndoMMLU and CN-Eval from effective dataset preparation, enhanced instruction-following via weight merging with minimal tuning, and resource efficiency through optimized TPU/GPU strategies for scalable training.
Future directions include:
  • Expanding to more underrepresented languages like Tagalog, Thai, and Vietnamese.
  • Developing enhanced evaluation frameworks with human judgment and multimodal input.
  • Applying the model to tasks such as translation, summarization, and content analytics.
MERaLiON demonstrates the potential for resource-efficient multilingual NLP, setting the foundation for broader linguistic coverage and consistent multilingual reasoning.
Acknowledgement
We extend our sincere gratitude to Nattadaporn Lertcheva, Xi Wang, Kui Wu, Yang Ding, Nabilah Binte Md Johan for their invaluable contributions to data, insightful discussions, and future work explorations.
This research is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority, Singapore under its National Large Language Models Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority, Singapore.
The resources and platforms provided by Singapore NSCC Aspire2A+ and The TPU Research Cloud. We thank all contributors and collaborators who have made this effort possible.